sanchez can do a certain job in 2 hours, whereas it takes barry 3 hours to do the same job. how long would it take them to do the job working together?
Answer by richwmiller(17219)

1/a+1/b=1/x
1/2+1/3=1/x
3/6+2/6=1/x
5/6=1/x
x=6/5
x=1.2 hours working together
check
x/a+x/b=1
x/2+x/3=1
1.2/2+1.2/3=1
0.6+0.4=1
1.0=1
ok
codewt